OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> 0.9° -9.1° 1.8° 11.8°

OpenStudy (anonymous):

i think its 0.9 but lol just wanna clarify

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

ok that was wrong sorry

OpenStudy (anonymous):

so let me see im confused on how to set it up lol would you mind showing me! the steps and stuff :) @misty1212

OpenStudy (anonymous):

lol i figured i did something wrong lol

OpenStudy (misty1212):

take the dot product, you get \(32\) if i am not mistake

OpenStudy (misty1212):

then divide by the norm of both vectors you know how to find that?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

not sure what the norm would be

OpenStudy (misty1212):

\[||<-5,-4>|=\sqrt{5^2+4^2}\] just like pythatoras

OpenStudy (anonymous):

mhm i see

OpenStudy (misty1212):

so \[\frac{32}{\sqrt{41}\sqrt{25}}\]

OpenStudy (misty1212):

then take the arc cosine

OpenStudy (anonymous):

mhm im watching

OpenStudy (misty1212):

now you need a calculator i would use this http://www.wolframalpha.com/input/?i=arccos%2832%2F%285*sqrt%2841%29%29%29

OpenStudy (anonymous):

so i put that in and i got really weird numbers lol

OpenStudy (anonymous):

0.0312398334 rad

OpenStudy (anonymous):

@misty1212 can you show me what it is please lol im confused

OpenStudy (misty1212):

your calculator is in radian mode, but your answers are in degrees

OpenStudy (misty1212):

either put your calculator in degree mode, or just click on the link i sent the answer is there, both in degrees and in radians

OpenStudy (anonymous):

hmm i just dont know what the answer would be feel like im mixing all my numbers up lol

OpenStudy (anonymous):

1.7899106065º i converted the radians to degrees

OpenStudy (anonymous):

oh its 1.8!!! okay